# Zero-Order Sharpness-Aware Minimization

### v1

@@ 1 Introduction
Zeroth-order (ZO) optimization has become indispensable in machine learning scenarios where gradient information is inaccessible or computationally prohibitive, such as black-box adversarial attacks (Ru et al., 2020; Hiranandani et al., 2021) and memory-constrained fine-tuning of large language models (LLMs) (Malladi et al., 2023a; Zhang et al., 2024b). Unlike first-order methods that rely on backpropagation, ZO algorithms estimate gradients solely through function evaluations, enabling applications in resource-limited environments (Liu et al., 2018b; Chen et al., 2019; Shu et al., 2024).
@@
However, traditional ZO methods suffer from high variance in gradient estimates, leading to slow convergence and suboptimal generalization, particularly in high-dimensional non-convex landscapes (Chen et al., 2019; Nazari et al., 2020). Adaptive ZO optimizers, such as ZO-AdaMM (Chen et al., 2019; Nazari et al., 2020), attempt to mitigate these issues by incorporating moment estimates for better scaling of updates. Yet, they underutilize historical information, resulting in noisy estimates and limited performance gains (Shu et al., 2025). Recent advances like R-AdaZO (Shu et al., 2025) refine moment utilization through variance reduction on first-moment estimates and improved second-moment approximations, achieving faster convergence. Meanwhile, sharpness-aware minimization (SAM) (Foret et al., 2021) has emerged as a powerful technique in first-order settings to enhance generalization by seeking flat minima, but its direct application to ZO is challenging due to the absence of gradients. SABO (Ye et al., 2024) extends SAM to zero-order optimization by reparameterizing objectives over Gaussian distributions and approximating sharpness-aware updates via stochastic gradients. Similarly, FZOO (Dang & Others, 2025) accelerates ZO fine-tuning of LLMs using batched one-sided estimates and adaptive step-sizes based on loss standard deviations, approaching Adam-like speeds with inference-level memory.
@@
Despite these progresses, existing ZO methods often trade off between memory efficiency, convergence speed, and generalization. For instance, MeZO (Malladi et al., 2023b) reduces memory to inference levels but requires significantly more forward passes than Adam. To bridge this gap, we introduce Zeroth-Order Sharpness-Aware (ZOSA) optimization, a novel adaptive ZO optimizer that integrates sharpness-aware mechanisms with refined variance reduction and adaptive scaling. ZOSA employs batched Rademacher perturbations for efficient one-sided gradient estimates, computes adaptive step-sizes using the standard deviation of batch losses, and incorporates a SAM-like perturbation at a scaled point to promote flat minima and better generalization. This design reduces variance in estimates (building on R-AdaZO (Shu et al., 2025)) while minimizing forward passes. Our contributions are summarized as follows:
@@
  • We propose ZOSA, an efficient zeroth-order (ZO) optimizer that seamlessly combines sharpness-aware updates with adaptive loss-variance-based scaling, with minimal inference-time memory overhead. The algorithm features a simple design, requiring no additional updates to a $\Sigma$ matrix (thus avoiding the extra storage burden and computational complexity associated with $\Sigma$ matrix updates), it offers fast computation, significantly fewer function queries, and accuracy superior to or on par with SABO, which is the state-of-the-art of shape-aware ZO algorithm, providing a practical and easy-to-implement solution for ZO optimization.
@@
  • We provide comprehensive theoretical analysis, including proofs of equivalence to a SAM rule, rigorous variance reduction bounds, and convergence guarantees under standard assumptions, further solidifying its theoretical foundations.
@@
  • Extensive experiments on synthetic problems and large language model (LLM) fine-tuning tasks demonstrate ZOSA’s superiority in convergence speed, resource efficiency, and performance stability, highlighting its potential in real-world applications.
@@ 2 Related Works
ZO optimization research primarily advances in gradient estimation and update rules, with growing emphasis on adaptivity, generalization, and applications to large-scale models like LLMs. Early ZO methods rely on finite-difference approximations for gradients, such as two-point estimates (Liu et al., 2018b; a). To address high-dimensional challenges, random direction sampling (e.g., Gaussian, Rademacher, or coordinate-wise) reduces query complexity while maintaining unbiased estimates (Chen et al., 2019; Shu et al., 2024; Nesterov & Spokoiny, 2011). Recent works like MeZO (Malladi et al., 2023b) apply these techniques to LLM fine-tuning, replacing backpropagation with forward passes to achieve inference-level memory usage. However, MeZO’s fixed step-sizes lead to slow convergence, often requiring 10-20× more iterations than first-order methods, as practitioners highlight its dependence on local Hessian rank rather than parameter count.
@@
ZO for LLM Fine-Tuning. With the rise of LLMs, ZO has been tailored for memory-efficient fine-tuning in resource-constrained settings. For instance, DP-ZO (Lin et al., 2024) introduces differential privacy into ZO for private LLM adaptation using forward-only perturbations. Quantized variants like QuZO (Qu et al., 2025) enable low-bit ZO fine-tuning, achieving performance comparable to MeZO on tasks like GLUE while reducing computational overhead. HiZOO (Cai et al., 2024) leverages diagonal Hessian approximations to enhance ZO updates, and FedMeZO (Zhang et al., 2024a) extends this to federated learning, proving convergence in distributed settings. SubZero (Chen et al., 2024a) tackles sparsity by optimizing in random subspaces, mitigating dimensionality curses in billion-parameter models. These methods collectively demonstrate ZO’s potential for LLM adaptation but often overlook generalization in non-convex landscapes (Zhang et al., 2024b; Malladi et al., 2023b).
@@
Adaptive ZO Optimizers and Variance Reduction. Adaptive methods improve upon basic SGD-like ZO by incorporating momentum and scaling. ZO-AdaMM (Chen et al., 2019; Nazari et al., 2020) adapts Adam’s moments to ZO but suffers from high-variance estimates in noisy environments. R-AdaZO (Shu et al., 2025) overcomes this by providing the first analysis of variance reduction via first-moment estimates and refining second moments for better geometry capture, yielding faster convergence than ZO-AdaMM. FZOO (Dang & Others, 2025) accelerates ZO through batched Rademacher perturbations and adaptive step-sizes based on batch loss standard deviations, emulating normalized-SGD (You et al., 2019) without momentum costs, and seamlessly integrates with PEFT techniques like LoRA (Hu et al., 2021) for further memory savings. Additional variance reduction approaches, such as LOZO (Chen et al., 2024b), use low-rank gradient estimators to capture low-dimensional structures in LLM loss landscapes (Zhou et al., 2025).
@@
Sharpness-Aware and Zero-Order Optimization. Sharpness-aware minimization (SAM) (Foret et al., 2021) promotes generalization in first-order optimization by seeking flat minima through neighborhood maximization, but it requires gradients. Extensions like VS-SAM (Liu et al., 2023) suppress variance in perturbations for stable training. In zero-order settings, SABO (Ye et al., 2024) reparameterizes objectives over Gaussian distributions to approximate sharpness-aware stochastic gradients, with proven convergence and generalization bounds. SharpZO (Yang et al., 2025) hybridizes this for vision-language models, enhancing ZO prompt optimization. These align with empirical findings that flat minima correlate with better performance (Dziugaite & Roy, 2017; Petzka & Sminchisescu, 2021; Andriushchenko & Flammarion, 2022). In LLM contexts, sharpness-aware ZO could alleviate overfitting in zero-order tuning (Sun et al., 2022b; 2023), though existing methods like SAM variants (Huang et al., 2024) are limited to gradient-based scenarios.
@@
ZOSA builds on these foundations by fusing variance-reduced adaptive scaling (Shu et al., 2025), efficient batched estimation, and sharpness-aware updates (Ye et al., 2024), while incorporating elements from recent LLM-specific ZO works (Lin et al., 2024; Cai et al., 2024), to provide a unified, memory-efficient framework for generalizable optimization.
@@ 3 Methodology
1:Model parameters $\theta$, loss function $L(\theta)$, sharpness-aware radius $\rho$, perturbation scale $\epsilon$, number of sampled directions $m$, learning rate $\eta$
@@
5:  Estimate gradient at the original point:
6:  Generate $m$ Rademacher perturbation vectors $u_{i}$ for $i=1$ to $m$
7:  Compute original loss $l_{0}=L(\theta)$
8:  for each $u_{i}$ do
9:   Perturb parameters: $\theta^{\prime}=\theta+\epsilon\cdot u_{i}$
10:   Compute perturbed loss $l_{i}=L(\theta^{\prime})$
11:   Restore parameters: $\theta^{\prime}=\theta$
12:  end for
13:  Estimate gradient $g_{t}=\frac{1}{m}\sum_{i=1}^{m}\frac{l_{i}-l_{0}}{\epsilon}\cdot u_{i}$
14:  Compute standard deviation for adaptive step size:
15:  $\sigma_{t}=\text{std}([l_{i}\text{ for }i\text{ in }1\text{ to }m])$
16:  Sharpness-aware perturbation:
@@
18:   Compute perturbation $\epsilon_{\text{sam}}=\rho\cdot\frac{g_{t}}{\sigma_{t}}$
@@
25:  Compute perturbed point loss $l_{\text{pert}}=L(\theta_{\text{pert}})$
@@
33:  $\sigma_{t,\text{pert}}=\text{std}([l_{i,\text{pert}}\text{ for }i\text{ in }1\text{ to }m])$
34:  Set adaptive learning rate $\eta_{\text{adaptive}}=\begin{cases}\frac{\eta}{\sigma_{t,\text{pert}}}&\text{if }\sigma_{t,\text{pert}}>0\\ \eta&\text{otherwise}\end{cases}$
35:  Restore original parameters:
36:  $\theta_{\text{pert}}=\theta$
37:  Update parameters:
38:  $\theta=\theta-\eta_{\text{adaptive}}\cdot g_{\text{pert}}$
39:end while
40:return $\theta$
@@ 3.1 Preliminaries
Consider a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}$, where a mini-batch $\mathcal{B}\subset\mathcal{D}$ has size $|\mathcal{B}|$. Let $\theta\in\mathbb{R}^{d}$ denote all trainable parameters of the LLM, and let $L(\theta;\mathcal{B})$ be the empirical loss on $\mathcal{B}$.
Classical ZO Gradient Estimation Given a perturbation radius $\epsilon>0$ and $z\in\mathbb{R}^{d}$ sampled as $z\sim\mathcal{N}(0,I_{d})$, the Classical ZO estimates the gradient on $\mathcal{B}$ via:
@@
From classical ZO to ZO-SGD. Replacing the back-propagation gradient in SGD with the classical ZO estimate gives the zeroth-order update $\theta_{t+1}=\theta_{t}-\eta_{t}\hat{\nabla}L(\theta_{t};\mathcal{B}_{t})$. MeZO realizes this update in-place with the memory tricks above and serves as a baseline for improvements. The introduction of Fast Zeroth-Order Optimizer (FZOO) is in Appendix A.1.
@@ 3.2 Motivation of ZOSA
To enhance generalization while maintaining efficiency, ZOSA builds upon FZOO by incorporating sharpness-aware minimization (SAM) principles. This integration targets flatter minima in the loss landscape. Specifically, ZOSA first estimates the gradient at the current parameters using batched Rademacher perturbations, computes an adaptive perturbation based on the estimated gradient normalized by loss variance, and then performs a second gradient estimation at the perturbed point for the final update. This dual-estimation approach combines FZOO’s memory-efficient zeroth-order strategy with SAM-like sharpness awareness, normalized via variance for stability in high-dimensional spaces.
Letting $u_{1},\dots,u_{m}$ be $m$ i.i.d. Rademacher random vectors in $\mathbb{R}^{d}$, we construct the initial gradient estimation at the original point $\theta_{t}$ by function value queries $l_{i}=L(\theta_{t}+\epsilon u_{i};\mathcal{B}_{t})$, and $l_{0}=L(\theta_{t};\mathcal{B}_{t})$. That is, the gradient estimate $g_{t}$ is computed by averaging $m$ one-sided difference estimates
$g_{t}=\frac{1}{\epsilon m}\sum_{i=1}^{m}(l_{i}-l_{0})u_{i}.$ (2)
@@
$\theta_{t+1}=\theta_{t}-\eta_{t}\frac{g_{\text{pert}}}{\sigma_{t,\text{pert}}},$ (6)
where $\eta_{t}$ is the base learning rate. The detailed implementation of ZOSA is outlined in Alg. 1.

### v2

@@ 1 Introduction
Zeroth-order (ZO) optimization has become indispensable in machine learning scenarios where gradient information is inaccessible or computationally prohibitive, such as black-box adversarial attacks (ru2020bayesopt; hiranandani2021blackbox) and memory-constrained fine-tuning of large language models (LLMs) (malladi2023; zhang2024b). Unlike first-order methods that rely on backpropagation, ZO algorithms estimate gradients solely through function evaluations, enabling applications in resource-limited environments (liu2018a; chen2019; shu2024).
@@
However, traditional ZO methods suffer from high variance in gradient estimates, leading to slow convergence and suboptimal generalization, particularly in high-dimensional non-convex landscapes (chen2019; nazari2020). Adaptive ZO optimizers, such as ZO-AdaMM (chen2019; nazari2020), attempt to mitigate these issues by incorporating moment estimates for better scaling of updates. Yet, they underutilize historical information, resulting in noisy estimates and limited performance gains (shu2025). Recent advances like R-AdaZO (shu2025) refine moment utilization through variance reduction on first-moment estimates and improved second-moment approximations, achieving faster convergence. Meanwhile, sharpness-aware minimization (SAM) (foret2021) has emerged as a powerful technique in first-order settings to enhance generalization by seeking flat minima, but its direct application to ZO is challenging due to the absence of gradients. SABO (ye2024) extends SAM to zero-order optimization by reparameterizing objectives over Gaussian distributions and approximating sharpness-aware updates via stochastic gradients. Similarly, FZOO (dang2025fzoo) accelerates ZO fine-tuning of LLMs using batched one-sided estimates and adaptive step-sizes based on loss standard deviations, approaching Adam-like speeds with inference-level memory.
@@
Despite these progresses, existing ZO methods often trade off between memory efficiency, convergence speed, and generalization. For instance, MeZO (malladi2023kernel) reduces memory to inference levels but requires significantly more forward passes than Adam. To bridge this gap, we introduce Zeroth-Order Sharpness-Aware (ZOSA) optimization, a novel adaptive ZO optimizer that integrates sharpness-aware mechanisms with refined variance reduction and adaptive scaling. ZOSA employs batched Rademacher perturbations for efficient one-sided gradient estimates, computes adaptive step-sizes using the standard deviation of batch losses, and incorporates a SAM-like perturbation at a scaled point (with effective radius $\rho/\epsilon$) to promote flat minima and better generalization. This design reduces variance in estimates (building on R-AdaZO (shu2025)) while minimizing forward passes. Our contributions are summarized as follows:
@@
  • We propose ZOSA, an efficient ZO optimizer that seamlessly combines sharpness-aware updates with adaptive loss-variance-based scaling, with minimal inference-time memory overhead. The algorithm features a simple design, requiring no additional updates to a $\Sigma$ matrix (thus avoiding the extra storage burden and computational complexity associated with $\Sigma$ matrix updates), it offers fast computation, significantly fewer function queries, and accuracy superior to or on par with SABO, which is the state-of-the-art sharpness-aware ZO algorithm, providing a practical and easy-to-implement solution for ZO optimization.
@@
  • We provide comprehensive theoretical analysis, including proofs of approximate equivalence to a normalized SAM rule with effective sharpness radius $\rho/\epsilon$, rigorous variance reduction bounds, and convergence guarantees under standard assumptions, further solidifying its theoretical foundations.
@@
  • Extensive experiments on synthetic problems and LLM fine-tuning tasks demonstrate ZOSA’s superiority in convergence speed, resource efficiency, and performance stability, highlighting its potential in real-world applications.
@@ 2 Related Works
ZO optimization research primarily advances in gradient estimation and update rules, with growing emphasis on adaptivity, generalization, and applications to large-scale models like LLMs. Early ZO methods rely on finite-difference approximations for gradients, such as two-point estimates (liu2018a; liu2018b). To address high-dimensional challenges, random direction sampling (e.g., Gaussian, Rademacher, or coordinate-wise) reduces query complexity while maintaining unbiased estimates (chen2019; shu2024; nes2017). Recent works like MeZO (malladi2023kernel) apply these techniques to LLM fine-tuning, replacing backpropagation with forward passes to achieve inference-level memory usage. However, MeZO’s fixed step-sizes lead to slow convergence, often requiring 10-20× more iterations than first-order methods, as practitioners highlight its dependence on local Hessian rank rather than parameter count.
@@
ZO for LLM Fine-Tuning. With the rise of LLMs, ZO has been tailored for memory-efficient fine-tuning in resource-constrained settings. For instance, DP-ZO (lin2024) introduces differential privacy into ZO for private LLM adaptation using forward-only perturbations. Quantized variants like QuZO (qu2025quzo) enable low-bit ZO fine-tuning, achieving performance comparable to MeZO on tasks like GLUE while reducing computational overhead. HiZOO (cai2024) leverages diagonal Hessian approximations to enhance ZO updates, and FedMeZO (zhang2024fedmezo) extends this to federated learning, proving convergence in distributed settings. SubZero (chen2024subzero) tackles sparsity by optimizing in random subspaces, mitigating dimensionality curses in billion-parameter models. These methods collectively demonstrate ZO’s potential for LLM adaptation but often overlook generalization in non-convex landscapes (zhang2024b; malladi2023).
@@
Adaptive ZO Optimizers and Variance Reduction. Adaptive methods improve upon basic SGD-like ZO by incorporating momentum and scaling. ZO-AdaMM (chen2019; nazari2020) adapts Adam’s moments to ZO but suffers from high-variance estimates in noisy environments. R-AdaZO (shu2025) overcomes this by providing the first analysis of variance reduction via first-moment estimates and refining second moments for better geometry capture, yielding faster convergence than ZO-AdaMM. FZOO (dang2025fzoo) accelerates ZO through batched Rademacher perturbations and adaptive step-sizes based on batch loss standard deviations, emulating normalized-SGD (you2019) without momentum costs, and seamlessly integrates with PEFT techniques like LoRA (hu2021lora) for further memory savings. Additional variance reduction approaches, such as LOZO (chen2024lozo), use low-rank gradient estimators to capture low-dimensional structures in LLM loss landscapes (zhou2025unraveling).
@@
Sharpness-Aware and Zero-Order Optimization. SAM (foret2021) promotes generalization in first-order optimization by seeking flat minima through neighborhood maximization, but it requires gradients. Extensions like VS-SAM (liu2023vssam) suppress variance in perturbations for stable training. In zero-order settings, SABO (ye2024) reparameterizes objectives over Gaussian distributions to approximate sharpness-aware stochastic gradients, with proven convergence and generalization bounds. SharpZO (yang2025sharpzo) hybridizes this for vision-language models, enhancing ZO prompt optimization. These align with empirical findings that flat minima correlate with better performance (dziugaite2017; petzka2021; andriushchenko2022). In LLM contexts, sharpness-aware ZO could alleviate overfitting in zero-order tuning (sun2022b; sun2023), though existing methods like SAM variants (huang2024crsam) are limited to gradient-based scenarios.
@@
ZOSA builds on these foundations by fusing variance-reduced adaptive scaling (shu2025), efficient batched estimation, and sharpness-aware updates (ye2024), while incorporating elements from recent LLM-specific ZO works (lin2024; cai2024), to provide a unified, memory-efficient framework for generalizable optimization.
@@ 3 Methodology
1:Model parameters $\theta$, loss function $L(\theta)$, downstream labeled dataset $\mathcal{D}$, batch size $|\mathcal{B}|$, sharpness-aware radius $\rho$, perturbation scale $\epsilon$, number of sampled directions $m$, learning rate $\eta$
@@
5:  Sample mini-batch $\mathcal{B}\subset\mathcal{D}$ uniformly at random
6:  Estimate gradient at the original point:
7:  Generate $m$ Rademacher perturbation vectors $u_{i}$ for $i=1$ to $m$
8:  Compute original loss $l_{0}=L(\theta;\mathcal{B})$
9:  for each $u_{i}$ do
10:   Perturb parameters: $\theta^{\prime}=\theta+\epsilon\cdot u_{i}$
11:   Compute perturbed loss $l_{i}=L(\theta^{\prime})$
12:   Restore parameters: $\theta^{\prime}=\theta$
13:  end for
14:  Estimate gradient $g_{t}=\frac{1}{m}\sum_{i=1}^{m}\frac{l_{i}-l_{0}}{\epsilon}\cdot u_{i}$
15:  Sharpness-aware perturbation:
16:  $\sigma_{t}=\text{std}([l_{0},l_{1},\dots,l_{m}])$
@@
18:   Compute perturbation $\epsilon_{\text{sam}}=\rho\cdot\frac{g_{t}}{\sigma_{t}+10^{-8}}$
@@
25:  Compute perturbed point loss $l_{\text{pert}}=L(\theta_{\text{pert}};\mathcal{B})$
@@
33:  $\sigma_{t,\text{pert}}=\text{std}([l_{\text{pert}},l_{1,\text{pert}},\dots,l_{m,\text{pert}}])$
34:  Set adaptive learning rate $\eta_{\text{adaptive}}=\begin{cases}\frac{\eta}{\sigma_{t,\text{pert}}+10^{-8}}&\text{if }\sigma_{t,\text{pert}}>0\\ \eta&\text{otherwise}\end{cases}$
35:  Update parameters:
36:  $\theta=\theta-\eta_{\text{adaptive}}\cdot g_{\text{pert}}$
37:  Restore parameters to original point: $\theta=\theta-\epsilon_{\text{sam}}$
38:end while
39:return $\theta$
@@ 3.1 Preliminaries
We consider the standard supervised fine-tuning (or prompt tuning) objective on a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}$: $L(\theta;\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\ell\big(h(\theta;x),y\big),$ where $\theta\in\mathbb{R}^{d}$ represents the trainable parameters, and $h(\cdot;x)$ denotes the frozen pretrained LLM with appended prompts.
Classical ZO Gradient Estimation Given a perturbation radius $\epsilon>0$ and $z\in\mathbb{R}^{d}$ sampled as $z\sim\mathcal{N}(0,I_{d})$, where $I_{d}\in\mathbb{R}^{d\times}$ is the identity matrix of dimension $d$, the Classical ZO estimates the gradient on $\mathcal{B}$ via:
@@
From Classical ZO to ZO-SGD. Let $\theta_{t}\in\mathbb{R}^{d}$ denote the trainable parameters at iteration $t$, $\mathcal{B}_{t}$ be the mini-batch sampled at iteration $t$, and $\hat{g}_{t}:=\hat{\nabla}L(\theta_{t};\mathcal{B}_{t})$ be the zeroth-order gradient estimator computed on $\mathcal{B}_{t}$ . Replacing the back-propagation gradient in SGD with this zeroth-order estimate directly yields the zeroth-order stochastic update
$\theta_{t+1}=\theta_{t}-\eta_{t}\hat{g}_{t},$
where $\eta_{t}>0$ is the learning rate. MeZO realizes this update in-place with the memory tricks above and serves as a baseline for improvements. The introduction of Fast Zeroth-Order Optimizer (FZOO) is in Appendix A.1.
@@ 3.2 Motivation of ZOSA
To enhance generalization while maintaining efficiency, ZOSA builds upon FZOO by incorporating SAM principles. This integration targets flatter minima in the loss landscape. Specifically, ZOSA first estimates the gradient at the current parameters using batched Rademacher perturbations, computes an adaptive perturbation based on the estimated gradient normalized by loss variance, and then performs a second gradient estimation at the perturbed point for the final update. This dual-estimation approach combines FZOO’s memory-efficient zeroth-order strategy with SAM-like sharpness awareness, normalized via variance for stability in high-dimensional spaces.
To obtain a low-variance zeroth-order gradient estimate, we draw $m$ independent Rademacher random vectors $u_{1},\dots,u_{m}\overset{\text{i.i.d.}}{\sim}\mathbb{R}^{d}$ (each component of $u_{i}\in\{-1,+1\}^{d}$ with equal probability $1/2$, independent across coordinates and samples). For a fixed perturbation radius $\epsilon>0$, we perform $m+1$ forward passes to query the following scalar function values:
$l_{0}$ $=L(\theta_{t};\mathcal{B}_{t}),$ (2)
$l_{i}$ $=L(\theta_{t}+\epsilon u_{i};\mathcal{B}_{t}),\quad i=1,\dots,m.$ (3)
The batched one-sided Rademacher gradient estimator at the original point $\theta_{t}$ is then constructed as
$\hat{g}_{t}=\frac{1}{m}\sum_{i=1}^{m}\frac{l_{i}-l_{0}}{\epsilon}u_{i}=\frac{1}{m\epsilon}\sum_{i=1}^{m}(l_{i}-l_{0})u_{i}.$ (4)
This estimator is unbiased for the directional derivative along each $u_{i}$ and exhibits significantly lower variance than single-sample ($m=1$) estimates.
@@
$\theta_{t+1}=\theta_{t}-\eta\frac{g_{\text{pert}}}{\sigma_{t,\text{pert}}},$ (8)
where $\eta$ is the base learning rate. The detailed implementation of ZOSA is outlined in Alg. 1.